OpenStudy (swag):

Which of the following sets of numbers is not a Pythagorean triplet? 12, 16, 20 14, 48, 50 16, 30, 34 20, 48, 56

OpenStudy (anonymous):

16,30,34 i think

OpenStudy (cwrw238):

check : is 34^2 = 16^2 + 30^2 ?

OpenStudy (cwrw238):

16^2 = 256 30^2 = 900

OpenStudy (agent0smith):

use \[\Large a^2 + b^2 = c^2\] where c is the biggest side length

OpenStudy (agent0smith):

When you find one where the left hand side doesn't equal the right, bingo.